OpenStudy (jackiem272):

Factor by grouping: 2-2sin^2(x)-3sinx=0

OpenStudy (solomonzelman):

2-2a^2-3a=0 I am saying "let sin(x)=a" so now..... -2a^2-3a+2=0 , multiply everything by -1, you get... 2a^2+3a-2=0 (2a-1)(a+2 )=0

OpenStudy (solomonzelman):

So, ( 2 Sin(x) - 1 ) ( Sin (x) + 2 ) = 0

OpenStudy (jackiem272):

Thank you :)

OpenStudy (solomonzelman):

np